### v2

@@ Abstract
Drug discovery typically consists of multiple steps, including identifying a target protein key to a disease’s etiology, validating that interacting with this target could prevent symptoms or cure the disease, discovering a small molecule or biologic therapeutic to interact with it, and optimizing the candidate molecule through a complex landscape of required properties. Drug discovery related tasks often involve prediction and generation while considering multiple entities that potentially interact, which poses a challenge for typical AI models. For this purpose we present MAMMAL - Molecular Aligned Multi-Modal Architecture and Language - a method that we applied to create a versatile multi-task multi-align foundation model that learns from large-scale biological datasets (2 billion samples) across diverse modalities, including proteins, small molecules, and genes. We introduce a prompt syntax that supports a wide range of classification, regression, and generation tasks. It allows combining different modalities and entity types as inputs and/or outputs. Our model handles combinations of tokens and scalars and enables the generation of small molecules and proteins, property prediction, and transcriptomic lab test predictions. We evaluated the model on 11 diverse downstream tasks spanning different steps within a typical drug discovery pipeline, where it reaches new SOTA in 9 tasks and is comparable to SOTA in 2 tasks. This performance is achieved while using a unified architecture serving all tasks, in contrast to the original SOTA performance achieved using tailored architectures.
The model code and pretrained weights are publicly available at https://github.com/BiomedSciAI/biomed-multi-alignment and https://huggingface.co/ibm/biomed.omics.bl.sm.ma-ted-458m.
@@
1 Introduction
@@
Drug discovery traditionally follows a multi-step pipeline that begins with identifying disease-associated proteins, progresses to finding compounds that can effectively target these proteins, and culminates in the optimization of drug candidates to meet rigorous standards for efficacy and safety. This process is both costly and labor-intensive [1], requiring extensive laboratory assays that measure drug-target interactions, assess cellular changes in disease-relevant cell lines, and validate therapeutic efficacy and safety [1, 2, 3]. Drugs in development can be either small molecules, which are stable, easy to manufacture, and suitable for oral delivery [4], or biologic therapeutics, such as engineered antibodies, which offer high specificity but require complex manufacturing and are typically administered by injection [5].
@@
Accelerating drug discovery has become a central focus in biomedical research, aiming to streamline target identification, drug design, and testing [6, 7, 8]. Analyzing gene expression profiles, particularly from single-cell RNA sequencing (RNA-seq), has emerged as a key tool for distinguishing between cell populations associated with different diseases [9, 10, 11, 12]. This analysis enhances our understanding of disease mechanisms, facilitates the identification of new drug targets, and allows for the examination of drug effects across various cell types [13, 14]. In drug design, trained generative models are utilized to synthesize new drug candidates for further exploration [15, 16, 17]. As high-throughput screening assays for measuring drug binding affinity are costly and challenging to scale, accurately predicting drug-target interactions can significantly enhance drug design, improving both efficacy and precision. Overall, predictive modeling of binding affinity, toxicity, and efficacy in the early stages of the pipeline can reduce reliance on expensive late-stage testing, ultimately saving time and resources in drug development.
When creating predictive and generative AI models, one of the key challenges in the field involves the question - how should different modalities and entities be combined as inputs/outputs to a model? [18] This uncertainty is especially dominant in predictive and generative tasks that involve interaction between entities, for example, predicting whether a specific antibody and a protein target are likely to bind or not.
In this work, we introduce the MAMMAL (Molecular Aligned Multi-Modal Architecture and Language) method, and develop a multi-align foundation model paired with a prompt syntax that integrates multiple data domains to support a wide range of drug discovery tasks. The multi-align model, which enables aligning multiple entities into a single prompt, has been extensively pre-trained on 2 billion samples from diverse datasets, using auxiliary tasks of mask infilling, denoising, generation, and classification. The multi-align model is compatible with both encoder-decoder and encoder-only architectures and effectively incorporates numerical values through continuous token embedding, enhancing numerical precision and reducing vocabulary size. We rigorously evaluate the multi-align model across 11 downstream tasks - spanning classification, regression, and generation - covering key stages of the drug discovery pipeline across three primary domains: small molecules, proteins, and gene expression profiles. The multi-align model achieves state-of-the-art performance in 9 tasks and matches top performance in the remaining 2 tasks. The model is publicly available on https://huggingface.co/ibm/biomed.omics.bl.sm.ma-ted-458m under the name ibm/biomed.omics.bl.sm.ma-ted-458m

### v3

@@ Abstract
Large language models applied to vast biological datasets have the potential to transform biology by uncovering disease mechanisms and accelerating drug development. However, current models are often siloed, trained separately on small-molecules, proteins, or transcriptomic data, limiting their ability to capture complex, multi-modal interactions. Effective drug discovery requires computational tools that integrate multiple biological entities while supporting prediction and generation, a challenge existing models struggle to address. For this purpose, we present MAMMAL - Molecular Aligned Multi-Modal Architecture and Language - a versatile method applied to create a multi-task foundation model that learns from large-scale biological datasets across diverse modalities, including proteins, small-molecules, and omics. MAMMAL’s structured prompt syntax supports classification, regression, and generation tasks while handling token and scalar inputs/outputs. Evaluated on eleven diverse downstream tasks, it reaches a new state of the art (SOTA) in nine tasks and is comparable to SOTA in two tasks, all within a unified architecture, unlike prior task-specific models. Additionally, we explored Alphafold 3 binding prediction capabilities on antibody-antigen and nanobody–antigen complexes showing significantly better classification performance of MAMMAL in 3 out of 4 targets.
@@
Figure 2: Comparison of antibody/nanobody–antigen binding prediction between MAMMAL and AlphaFold 3 (AF3). (a) HER2 extracellular domain (ECD) binder/non-binder discrimination scores for MAMMAL versus AF3. (b-c) Binding prediction performance for nanobodies (VHHs) against three antigens: CD206, VWF, and TBG (green = binders, red = non-binders; mean ± SD; unpaired two-sided Student’s t-test; *$P<0.05$, ****$P<0.0001$). (b) MAMMAL evaluation. (c) AF3 ipTM (Interface Predicted TM-score) evaluation. (d) HER2 ECD structure with representative binder/non-binder predictions. Therapeutic antibodies Trastuzumab (blue) and Pertuzumab (purple) shown for comparison. AF3 predicts both binders and non-binders interacting with the same domain, distinct from known therapeutic epitopes. (e) TBG structure with binding/non-binding VHHs. AF3 predicts distinct poses for binders vs non-binders.
@@
Despite major advances in understanding health and disease, many complex biological processes remain poorly understood. Recent advances in artificial intelligence have given hope that longstanding health challenges may prove treatable with AI-based methods  [1] and that the high failure rate of drugs in clinical trials, where approximately $90\%$ of drug candidates do not achieve approval [2], can be improved. Over the past century, the process of translating biological discoveries into effective treatments has matured. The drug discovery journey follows a multi-step pipeline that begins with identifying disease-associated proteins, progresses to finding compounds that can effectively target these proteins, and culminates in optimizing drug candidates to meet rigorous standards for efficacy and safety. This process is both costly and labor-intensive, requiring extensive laboratory assays that measure drug-target interactions, assess cellular changes in disease-relevant cell lines, and validate therapeutic efficacy and safety [3, 4, 5]. Drugs in development can be either small molecules, which are stable, easy to manufacture, and suitable for oral delivery [6], or biologic therapeutics, such as engineered antibodies, which offer high specificity but require complex manufacturing and are typically administered by injection [7].
@@
Accelerating drug discovery and improving the accuracy of its steps have become a central focus in biomedical research, with the goal of streamlining target identification, drug design, and testing [8, 9, 10]. Analyzing gene expression profiles, particularly from single-cell RNA sequencing (scRNA-seq), has emerged as a key tool for distinguishing between cell populations associated with different diseases [11, 12, 13, 14]. This analysis enhances our understanding of disease mechanisms, facilitates the identification of new drug targets, and allows for the examination of drug effects across various cell types [15, 16]. In drug design, trained generative models are utilized to synthesize new drug candidates for further exploration [17, 18, 19]. As high-throughput screening assays for measuring drug binding affinity are costly and challenging to scale, accurately predicting drug-target interactions can significantly enhance drug design, improving both efficacy and precision. Overall, predictive modeling of binding affinity, toxicity, and efficacy in the early stages of the pipeline can reduce reliance on expensive late-stage testing, ultimately saving time and resources in drug development.
When creating predictive and generative AI models, one of the key challenges in the field is determining how different modalities and entities should be combined as inputs and outputs to a model [20]. This presents a significant hurdle in predictive and generative tasks that involve interactions between entities, for example, predicting whether a specific small-molecule drug and a protein target are likely to bind. While the representation of small molecules and proteins as sequences, such as SMILES and amino acid chains, has demonstrated its value over the years, the approach and benefits of representing transcriptomic data as a sequence have only recently been studied, leading to multiple proposals [21].
In this paper, we propose a cross-domain model that leverages sequence representation for three modalities: small molecules, proteins, and transcriptomic data. Specifically, we introduce MAMMAL (Molecular Aligned Multi-Modal Architecture and Language), a foundational model designed to address the challenges associated with drug discovery tasks. Our key contributions are: (1) a novel multi-align method that combines data from small molecules, proteins, and gene expression using a flexible, structured prompt syntax. This approach enables the model to handle the complex, multi-entity interactions critical for drug discovery tasks such as binding prediction; (2) an architecture specifically designed for drug discovery, featuring seamless integration of numerical values directly into its embedding space via a projection layer. This improves precision and avoids the limitations of alternative methods, making it ideal for tasks such as predicting binding affinities or the integration of a 3D structure. Additionally, the model’s hybrid design optimizes performance across classification, regression, and generation tasks, all essential in different stages of drug development; and (3) A pre-trained model trained on over 2 billion samples across tasks like spans infilling, denoising, and binding prediction. By achieving SOTA results on 9 out of 11 benchmarks, MAMMAL demonstrates its ability to streamline key steps in the drug discovery pipeline. The model and pre-trained weights are publicly available on https://huggingface.co/ibm/biomed.omics.bl.sm.ma-ted-458m under the name ibm/biomed.omics.bl.sm.ma-ted-458m, providing a platform for advancing research and innovation in biomedicine.
1 Results
To evaluate the performance and generalization capabilities of ibm/biomed.omics.bl.sm.ma-ted-458m, we selected a diverse set of existing benchmarks spanning multiple task types and stages of the drug discovery pipeline, prioritizing benchmarks with clearly defined splits when those were available. We assessed model quality through a fine-tuning-based evaluation strategy, where the pretrained model is adapted to each benchmark and compared against specialized state-of-the-art (SOTA) models. The evaluation methodology and fine-tuning protocol as well as detailed descriptions of each benchmark—including background, significance for drug discovery, prior models, and data statistics—are provided in the subsections below. A summary of performance across tasks is presented in Table 1 and visualized in Figure 1(E), and representative encoder-decoder examples are shown in Table LABEL:table_task_queries.
AlphaFold [22], whose development contributed to the 2024 Nobel Prize in Chemistry, revolutionized protein structure prediction. Its extension AlphaFold-Multimer [23] enabled modeling of antibody-antigen complexes, while AlphaFold 3 (AF3) [24] further improved accuracy and added nucleic acid/small molecule support. Motivated by AF3’s reported advances, we evaluated its performance on therapeutic Antibody and Nanobody complexes (Subsection 1.10). Comparative analysis showed MAMMAL achieved better classification performance than AF3 in three of four targets (Figure 2).
Table 1: Comparison of SOTA and MAMMAL Performance Across Benchmarks
Benchmark Domain Type Metric SOTA MAMMAL Imp.
Cell type GE cls $\uparrow$ F1 [25] 0.710 0.763±0.012 7.5 %
BBBP SM cls $\uparrow$ AUROC [26] 0.937 0.957±0.006 2.2 %
ClinTox SM cls $\uparrow$ AUROC [26] 0.948 0.986±0.007 4.0 %
DTI Prot.+SM reg $\downarrow$ NRMSE 0.942±0.028 [33] 0.906±0.011 3.8 %
00footnotetext: For NRMSE, lower is better. For other metrics (AUROC, CDRH3-AAR, Pearson, Spearman, and F1), higher is better. Each row shows the results from a MAMMAL model fine-tuned from ibm/biomed.omics.bl.sm.ma-ted-458m for the corresponding task. Abbreviations: In the ”Type” column: ”cls” = classification; ”reg” = regression; ”gen” = generation; ”Imp.” = improvement (percentage) of our model over SOTA. In the ”Domain” column: ”GE” = genes expression; ”SM” = small molecule; ”Prot.” = protein.
1.1 Evaluation
1.2 Cell Type Annotation
Cell type prediction enables researchers to distinguish between different cell populations, such as those associated with various diseases [11, 12, 13, 14]. It is also crucial for understanding how diseases or drugs affect different cell types. In recent years, a variety of methods have been developed for this task, including approaches based on marker genes, correlation-based techniques, and annotation using classification [34]. Recent advances in transformer-based and large-scale foundation models [35, 25, 36] have shown improved performance.
The input for this task is single-cell gene expression data. The benchmark we used is based on the Zheng68k dataset [37], which is composed of human peripheral blood mononuclear cells and is widely used for evaluating cell-type annotation performance, due to the similarity of the cell types involved. The dataset contains 68,579 cells across 11 cell types and originally included 32,738 genes, which after removing non-expressed genes leaves 20,387 genes in the benchmark. Preprocessing involved normalization, log transformation of expression values, followed by binning. Similar to the approach in [38], our model uses a ranked list of expressed gene names, ordered by their expression levels, as input. The label to predict is provided in the cell ontology format “CL:NNNNNN” (see Table LABEL:table_task_queries).
1.3 BBBP and ClinTox
To ensure the development of safe and effective drugs, candidates must satisfy rigorous criteria related to both efficacy and safety. In this study, we selected two relevant benchmarks from MoleculeNet [39], a widely used suite of benchmarks for evaluating machine learning models on small-molecule drug properties: BBBP and ClinTox. The BBBP benchmark focuses on predicting the ability of drugs to penetrate the blood-brain barrier, a critical consideration for drugs targeting the central nervous system. The ClinTox benchmark comprises two related tasks: (1) predicting failure in clinical toxicity trials, and (2) predicting FDA approval status. The overall performance on ClinTox is reported as the average performance across these two tasks.
MoLFormer [26], a well-established model for molecular embeddings trained on 1.1 billion SMILES sequences, has achieved state-of-the-art performance on both the BBBP and ClinTox benchmarks. In our study, we adopted the benchmarks from [26], which provided predefined splits for training, validation, and testing. As shown in Table  1, MAMMAL surpassed MoLFormer on both benchmarks, achieving an average area under the receiver operating characteristic curve (AUROC) score of 0.957 on BBBP and 0.986 on ClinTox, representing improvements of 2.2% and 4%, respectively, over the state of the art.
1.4 Cancer-Drug Response
For our study, we used three subsets of the GDSC database: GDSC1 and GDSC2, available through the Therapeutics Data Commons (TDC) [44], and referred to in the paper as Cancer-Drug Response 1 and Cancer-Drug Response 2, respectively; and a subset published in [28], referred to as Cancer-Drug Response 3. Table S3 summarizes the number of cell lines, drugs, and cell-drug pairs in these datasets. We used the random splits provided by TDC for Cancer-Drug Response 1 and 2, while for Cancer-Drug Response 3, we followed the split methodology outlined in [28], reserving 5% of the data for the test set, stratified by TCGA [45] pathways associated with the cancer cell lines.
During fine-tuning, we used only gene-expression profiles and SMILES representations of drugs, as shown in the example prompt in Table LABEL:table_task_queries. Similar to the input format for cell type annotation, gene-expression profiles were provided as ranked lists of gene names based on their expression levels. For predicting continuous IC50 values, MAMMAL was utilized in regression mode, taking advantage of its built-in support for floating-point scalar predictions. As demonstrated in Table 1, our model outperforms the current SOTA models for Cancer-Drug Response 1 and 2, achieving a 3.4% increase in Pearson correlation values. Additionally, it yields results comparable to the SOTA for the Cancer-Drug Response 3 benchmark, with a slight improvement of 0.5%.
1.5 Targeted Antibody Design
Antibodies are a family of proteins produced by the immune system to neutralize foreign antigens and are of particular interest due to their high specificity and strong binding to target molecules [46, 47]. These characteristics have made them a crucial class of therapeutics, driving significant research efforts into the design of new antibody-based drug candidates [48, 49, 50, 7]. Antigen-binding fragments (Fabs) are the antibody fragments that bind to antigens. It is composed of one constant domain and one variable domain of each of the heavy and light chains. Each variable region is further divided into four framework (FR) regions and three complementarity-determining regions (CDRs). While FR regions are typically conserved, CDRs exhibit significant variation in their amino acid composition and are generally the primary determinants of binding affinity to the target antigen. When designing novel antibodies for a specific antigen, the typical approach is to explore alternative CDRs that could produce a new, functional antibody with high binding affinity to the target [46, 47, 51, 29].
Recently, several deep learning methods have been developed for targeted antibody design, framing CDR prediction as an infilling task [51, 52, 53, 54, 55, 29, 56]. These models predict missing CDR regions, represented by MASK tokens, using the amino acid sequences of both the antigen and the antibody’s FR regions. While prior approaches often relied on structural data, this information is scarce and challenging to obtain [57]. In contrast, we fine-tune MAMMAL for the targeted antibody design task using only the sequence data of the antigen and the sequence of the antibody’s FR regions.
1.6 T-Cell Receptor-Epitope Binding
T-cell receptor (TCR) binding to immunogenic peptides (epitopes) presented by major histocompatibility complex molecules is a critical mechanism in the adaptive immune system, essential for antigen recognition and triggering immune responses. The TCR repertoire exhibits considerable diversity, consisting of an $\alpha$-chain and a $\beta$-chain that function together to enable T cells to recognize a wide array of epitopes. The $\beta$-chain is especially significant, as it is crucial for the early stages of T-cell development and possesses greater variability, which enhances the TCR’s capacity to identify diverse pathogens effectively. However, understanding the specific interactions between TCRs and epitopes remains a significant challenge due to the vast variability in TCR sequences. Accurate prediction of TCR-peptide binding from sequence data could revolutionize immunology by offering deeper insights into a patient’s immune status and disease history. This capability holds potential applications in personalized immunotherapy, early diagnosis, and the treatment of diseases such as cancer and autoimmune disorders. In silico tools designed to model TCR-peptide interactions could also facilitate the study of therapeutic T-cell efficacy and assess cross-reactivity risks, presenting a transformative opportunity for precision medicine.
We evaluated the model on the task of predicting TCR-epitope binding from sequence data using the Weber benchmark ([31], [58]), which consists of 47,182 TCR $\beta$-chain  epitope pairs. This dataset covers 192 distinct epitopes and includes 23,139 unique TCR $\beta$-chain sequences, with 50% of the pairs serving as negative samples created by randomly pairing TCR sequences with epitopes they are not known to bind with. The dataset also includes the CDR3 subsequence for each TCR $\beta$-chain , the most hypervariable region of the chain. We used 10-fold cross-validation. The folds were pre-defined in [31]. Fine-tuning involved three concurrent tasks: TCR $\beta$-chain mask infilling and two classification tasks: (i) TCR $\beta$-chain epitope binding prediction and (ii) TCR $\beta$-chain -CDR3 epitope binding prediction. Here, we report the performance only for the TCR $\beta$-chain epitope binding prediction task. As depicted in Table 1, our model achieved an average AUROC of 0.879, representing a statistically significant improvement of 2% over the SOTA, as our result falls outside the SOTA’s confidence interval.
1.7 Protein-Protein Interaction - $\Delta$$\Delta$G Prediction
$\Delta G=kT\ln(K_{D}),$ (1)
where $k$ is the Boltzmann constant and $T$ is the temperature [59].
$\Delta\Delta G=\Delta G_{\text{mutant}}-\Delta G_{\text{wild-type}}.$
The SKEMPI dataset [59] provides information on changes in thermodynamic parameters, including $\Delta G$, and kinetic rate constants due to mutations in protein-protein (PP) complexes whose structures are available in the Protein Data Bank [63]. This dataset is extensively utilized in the literature for predicting the effects of mutations on binding affinity, particularly in the context of $\Delta\Delta G$. A subset of SKEMPI comprising 1,131 samples of single mutations, S1131, is widely used. We adopt S1131 as our benchmark for predicting protein-protein $\Delta\Delta G$ and follow the common practice of reporting 10-fold cross-validation performance on this subset. The input prompt for our model includes the reference “wild-type” version of the complex and the corresponding mutated version, comprising only sequence data without any structural information. We leverage MAMMAL’s support for floating-point scalars to predict continuous $\Delta\Delta G$ in a regression task setting. Performance results are presented in Table 1. As shown in Table 1, our model achieved an average Pearson correlation of 0.852, significantly exceeding the previous sequence-only SOTA of 0.663. Compared to models that incorporate structural data, our model’s performance remains competitive, falling just 1.6% short of the SOTA performance of 0.866 [60].
1.8 Drug-Target Interaction
1.9 Antibody-Antigen Binding Prediction
The HER2 dataset was divided into train (70%), validation (15%) and test (15%) sets.
Finetuning involved feeding the target antigen sequence as well as the entire heavy-chain variable region as input and predicting binding to the target sequence. As depicted in Table 1, our model achieved an average ROCAUC of 0.928, slightly surpassing the SOTA, which incorporated structural data, unlike our model.
1.10 Comparison of AlphaFold 3 and MAMMAL in Predicting Antibody–Antigen and Nanobody–Antigen Binding
Accurate prediction of antibody–antigen and nanobody–antigen interactions is essential for evaluating therapeutic efficacy and guiding protein engineering. Although AlphaFold 3 (AF3) is not explicitly designed as a binary protein–protein interaction (PPI) classifier, it can infer binding likelihood when the predicted PPI structure exhibits a well-packed interface with high confidence, as measured by predicted template modeling (pTM) or interface predicted template modeling (ipTM) scores. Recent studies suggest that these scores correlate with true binding events [66, 67]. We conducted an exploratory comparison of AF3’s ability to distinguish binders from non-binders against a fine-tuned MAMMAL model.
We first evaluated the extracellular domain (ECD) of HER2, a well-characterized therapeutic antigen with known binding epitopes. Figure 2a compares the score distributions for MAMMAL and AF3 in classifying binders versus non-binders. Due to the computational demands and limited availability of AF3, we downsampled the test set to 60 examples (30 positive and 30 negative). We used the HER2-specific MAMMAL model described in Subsection 1.9, which achieved strong discriminative performance (AUROC = $0.93$; $P<0.0001$, unpaired two-sided $t$-test). Conversely, AF3 exhibited no meaningful separation (AUROC $\approx 0.5$) across all tested configurations. Furthermore, AF3’s predicted ECD binding sites differed from the known epitopes of the FDA-approved antibodies trastuzumab and pertuzumab, as visualized in Figure 2d. Extended analyses are provided in Appendix E.
We next evaluated three structurally diverse targets: (i) CD206 (a 1500-amino-acid transmembrane protein), (ii) von Willebrand Factor (VWF; a $>$2000-amino-acid serum protein), and (iii) thyroxine-binding globulin (TBG; a 384-amino-acid serum protein). Binders were nanobodies sourced from SAbDab-nano [68], patents, and proprietary datasets, while non-binders comprised random nanobodies against unrelated targets and non-binders from a phage display library. From the full dataset (154 sequences: 46 binders and 108 non-binders), we selected 50% (binders and non-binders) for MAMMAL finetuning and 50% for evaluation. The fine-tuned MAMMAL model exhibited strong performance on CD206 and VWF (AUROC $=0.995$) and $0.83$ respectively) but performed poorly on TBG (AUROC = $0.7$), as shown in Figure 2b. In contrast, AF3 showed the opposite trend, achieving strong performance on the smaller TBG target (AUROC = $0.98$) but weak discrimination for the larger CD206 and VWF proteins (AUROC $<0.57$), as shown in Figure 2c. Structural analyses revealed distinct AF3-predicted binding sites for binders versus non-binders (Figure 2e). Extended analyses for all targets are provided in Appendix E.
2 Discussion
The rapid advancement of biomedical AI foundation models has focused on different aspects of cellular content. These include RNA expression-based models (e.g., [35, 36, 38]), protein-based models (e.g., [69, 70]), and models of small molecules interacting with proteins (e.g., [26]). As a result, we are moving closer to uncovering previously unknown biological processes. We may even be on the brink of solving the fundamental challenge of building a virtual cell. This concept has recently been proposed as a futuristic AI-based tool for novel biological discoveries and the design of new treatments [71, 72]. A crucial first step toward this goal is developing methods to integrate these diverse models. The MAMMAL approach proposed in this paper enables such integration. It reformulates tasks as sequence-to-sequence problems and introduces several key architectural enhancements: support for both encoder-only and encoder-decoder modes, a multi-domain extensible syntax for inputs and outputs, and the direct handling of numerical values through continuous token embeddings. A model was pretrained using this approach while aligning inputs across diverse datasets — including small molecule, protein, antibody, and gene expression data — using a variety of pretraining tasks. This multi-align approach enables the integration of cross-domain pharmaceutical knowledge into a single model and facilitates effective transfer learning to fine-tuned models for downstream applications. Demonstrated state-of-the-art performance of the multi-align fine-tuned models across diverse tasks, spanning multiple data domains and stages of the drug discovery pipeline, firmly establishes the MAMMAL approach as a powerful tool with transformative potential.
We also explored AF3 capabilities in predicting binding. We observed underwhelming performance of AF3 in three out of four cases, which might be a result of a built-in bias in the AF3 training process, such as exposing the model only to true positive binders. Another hypothesis is that sequence-based models are capable of learning from a more varied collection of protein chains, including proteins with Intrinsically Disordered Regions (IDR). Additionally, AF3 was not directly trained on separating binders and non-binders, which can also explain its underwhelming performance on this task. Going forward, we plan to expand MAMMAL to natively support 3D input and output, to enjoy the benefits of both sequence and structure-based methods.
In recent months, the world has rapidly transitioned to embrace agentic workflows, an advanced methodology that leverages large language models (LLMs) and specialized agents to automate and enhance discovery processes traditionally performed by humans [73, 74]. This shift marks a significant evolution in how research and development are conducted across industries. The fine-tuned models derived from MAMMAL seamlessly integrate into this paradigm, offering highly specialized capabilities that enhance efficiency and accuracy. By capitalizing on the power of multi-domain learning and tailored fine-tuning, these models are poised to revolutionize workflows, breaking new ground and unlocking unprecedented potential in the field of biomedical discovery.